OpenStudy (anonymous):

Two forces are acting upon an object of mass 20kg. One force has a magnitude of 50N and is directed toward the positive x axis. The second force has a magnitude of 25N and is directed in the positive y axis. What is the net force on the object? What angle does the net force make with the x axis? What is the acceleration of the object in this direction? Explain.

OpenStudy (apoorvk):

 now the resultant force is.\[\sqrt{ Fx^{2} + Fy^{2}}\] so net force is sq root of (3025) = 55 N F(net)=ma therefore, a=F/m = 2.75 m/sec in the direction of the net force.